Log10+Log10(x+2)-Log10(x-1)=2
Answer by fractalier(6550)

Let's rewrite what it is given as
Log10+Log10(x+2)-Log10(x-1)=2
1 + log(x+2) - log(x-1) = 2
log [(x+2)/(x-1)] = 1
(x+2)/(x-1) = 10
x+2 = 10(x-1)
x + 2 = 10x - 10
12 = 9x
x = 12/9 = 4/3